# MBIAS: Mitigating Bias in Large Language Models While Retaining Context

### v2

@@ Abstract
In addressing the critical need for safety in Large Language Models (LLMs), it is crucial to ensure that the outputs are not only safe but also retain their contextual accuracy. Many existing LLMs are safe fine-tuned either with safety demonstrations, or rely only on adversarial testing. While able to get safe outputs, they often risk losing contextual meaning as they mitigate bias and toxicity. In response, we present MBIAS, a LLM framework instruction fine-tuned on a custom dataset specifically designed for safety interventions. MBIAS aims to address the significant issues of bias and toxicity in LLMs generations that typically manifest as underrepresentation or negative portrayals across various demographics, including inappropriate linguistic mentions and biased content in social media. We experiment on MBIAS for safety interventions using various configurations, and demonstrate more than a 30% reduction in overall bias and toxicity while successfully retaining key information. Additionally, a demographic analysis on an out-of-distribution test set confirms the robustness of our approach, with reductions in bias and toxicity exceeding 90% across various demographics. The dataset and instruction fine-tuned MBIAS are made available to the research community  [^1] https://huggingface.co/newsmediabias/MBIAS.
@@
Shaina Raza1thanks: Corresponding author: shaina.raza@vectorinstitute.ai, Ananya Raval1, Veronica Chatrath1 shaina.raza, ananya.raval,veronica.chatrath@vectorinstitute.ai 1Vector Institute for Artificial Intelligence
@@ 1 Introduction
The rise of generative artificial intelligence (AI) has brought forth substantial ethical and safety challenges, raising concerns of misuse, misinformation, and bias of the generated information Wach et al. (2023). Recent initiatives aim at aligning large language models (LLMs) with ethical norms. These efforts prioritize mitigating biases and enhancing values such as inclusivity, fairness, and non-discrimination Guo et al. (2023). To address harmful, biased, or misleading content (referred to as ‘biased texts’ herein), foundational strategies include implementing safety guardrails in the outputs generated by LLMs. These guardrails provide guidelines and boundaries to ensure AI applications align with ethical standards and societal expectations Att (2024).
@@
Methods such as red-teaming Perez et al. (2022), human feedback during pre-training Casper et al. (2023), and data augmentation methods (e.g., balanced sampling, paraphrasing, or counterfactual data generation) Sadasivan et al. (2023) are often used to reduce biases in LLMs, with the goal of making them safer and more aligned with human values Ganguli et al. (2022); Korbak et al. (2023). In the fine-tuning phase, techniques like instruction tuning, reinforcement learning from human feedback (RLHF), and safety context distillation Ouyang et al. (2022); Qi et al. (2023); Bai et al. (2022) are also used to address unsafe behaviors and improve the model’s ability to generalize.
@@
Despite advancements in the implementation of LLM safety measures, one concern is the loss of actual context or meaning of the original text Schlicht et al. (2024). This delicate balance between two competing goals — reducing biases in the text and preserving informational integrity — highlights a central paradox in bias reduction efforts. Catastrophic forgetting, which occurs when a model forgets previously learned information while acquiring new knowledge, is particularly an issue worth noting while implementing safety mechanisms Luo et al. (2023). Demonstrating the understanding of language models post-safety interventions has thus become a topic of research and discussion Nadeem et al. (2021); Schick et al. (2021).
@@
Recent studies indicate that incorporating even a modest number of safety-focused examples during the instruction-tuning stage can effectively mitigate certain safety concerns Bianchi et al. (2023); Inan et al. (2023); Bai et al. (2022). State-of-the-art LLMs such as GPT-4, Mistral, PaLM, LLaMA-2 and 3, and Claude-2 have been further fine-tuned using high-quality demonstrations, including perturbations and adversarial prompts, to enhance safety. While not enirely foolproof, this safety-tuning enables LLMs to reduce biases in their outputs. Advancing beyond mere demonstrations, we propose instruction fine-tuning a LLM on safety mechanisms. This approach aligns with prior research that emphasizes the importance of the quality and breadth of instruction-tuning data for developing proficient and reliable instruction-following models Touvron et al. (2023); Wang et al. (2024).
@@
  • We curate an instruction-tuning dataset with a focus on safety considerations. This dataset comprises paired examples: one containing potentially unsafe elements, such as stereotypes or prejudices, and its corresponding benign (safe) counterpart. The dataset was carefully annotated by human annotators.
  • We propose MBIAS (Mitigating Bias through LLM), an instruction-fine tuned model, built on Mistral2-7B-instruct Jiang et al. (2023). The fine-tuning process involves training the model with our custom dataset that contains examples of both unsafe and safe instances, guiding the model to recognize biases and generate safe responses that retain the meaning of the original input.
  • MBIAS can be adapted for use with other LLMs. For simplicity and ease of implementation, we employ parameter-efficient fine-tuning Ding et al. (2023) to train the model. Similar to Llama Guard Inan et al. (2023), we aim to release model weights with the smallest available model to enhance user-friendliness.
@@ 2 Related Works
Safety in LLMs Establishing safety measures and protocols is crucial to upholding trust in generative AI and LLMs. Many methods have been proposed to address specific biases in language models. Traditional methods to ensure safe outputs includes examining the embedding space of words and sentences to mitigate biases in the text. Embedding-space-based methods are often applied after training, requiring little-to-no fine-tuning of the model. These methods function as post-processing debiasing steps Liang et al. (2020a); Ungless et al. (2022). Subtraction-based methods are also used to remove biases in language models, which map the embedding space into a neutral one Bolukbasi et al. (2016), maintaining equal distance between non-gendered words and pairs of gendered words. Another method is to ensure that gender-neutral words are orthogonal to the gender direction Zhao et al. (2018). In a related work Zhao et al. (2019), the gendered words are replaced with their opposites in the original training data, and the model is trained on both the original and augmented data. Other methods include efficient fine-tuning for debiasing Gira et al. (2022) and fine-tuning only a small portion of parameters on debiasing data Gira et al. (2022).
@@
Debiasing practicality and reliability also depend on dataset selection, evaluation methods, and demographic coverage. Datasets like RedditBias Barikeri et al. (2021), WinoBias Zhao et al. (2019), HolisticBias Smith et al. (2022), RealToxicityPrompts Gehman et al. (2020), and others discussed in Chang et al. (2024) are frequently used for evaluating bias reduction in models. A variety of bias evaluation metrics are used, such as WEAT Liang et al. (2020b), Perspective API API (2024), StereoSet Stereotype Score Nadeem et al. (2021), and LLM alignment methods Wang et al. (2023); Chang et al. (2024).
@@
Numerous considerations on their adoption and integration into production settings are currently under observation. Our focus lies in safety implications for instruction-tuned models: specifically, how these models respond safely to biased instructions, while retaining context. Recent research indicates that training LLMs on demographically perturbed data results in fairer models Qian et al. (2022). We explore whether training on unsafe-benign text pairs achieves better fairness in LLMs, while retaining knowledge.
@@ 3.1 Dataset Preparation
The dataset for this study, extracted from our vast collection of about 3.7M records [^2] Anonymous due to double blind review., spans diverse content from news and social media platforms. This dataset is entirely in English and covers 200+ safety risk aspects. We chose a statistically significant subset of 8.5k records for this work, balancing diversity with computational efficiency, ensuring representation across various aspects of bias in news media.
Data Labeling and Annotation Procedure: During labeling, each instance of biased text (unsafe) is transformed into its benign, safe, and debiased counterpart. In the initial step, we leverage GPT-4 from OpenAI OpenAI to generate the benign versions of the data. An illustrative prompt (without demonstration) is presented below, while demonstrations are provided in Appendix A:
@@
The dataset schema  [^3] We make our dataset available under the https://creativecommons.org/licenses/by-nc/4.0/ licence. is given in Listing 1 below:
@@
To ensure seamless interoperability and integration across computational environments, our dataset follows the Alpaca dataset format Dubois et al. (2023). This structure comprises an ‘instructions’, ‘input’, and ‘output’.
@@ 3.2 Model Training
To develop MBIAS, we employ QLoRA (Quantized Language Model Optimization with Low Rank Adapters) Belkada et al. (2023), a Parameter-Efficient Fine-tuning (PEFT) technique using bitsandbytes, alongside the HuggingFace transformers Trainer, to fine-tune the Mistral-7B-instruct-0.2 model with our specialized instruction dataset. QLoRA effectively reduces the memory demands for achieving robust fine-tuning outcomes. It balances accuracy with resource efficiency through a 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers.
@@ 4.1 Experimental Setting
Table 2: Hyperparameters used
@@ 4.2 Evaluation Data, Metrics, and Baselines
To evaluate the level of bias and toxicity before and after implementing safety interventions using our methodology, we utilized a mix of scoring methods: LLM-based scoring, statistical measures, such as true positives rate, and qualitative measures. The LLM-based scoring metrics used in this study, through DeepEval AI (2024), are defined as follows:
@@
The Knowledge Retention metric measures whether the LLM retains factual information from the input in its generated output. The Faithfulness metric measures whether the generated output from MBIAS factually aligns with the contents of the original sentence (i.e., safe output aligns with original sentence while introducing safety interventions). The Answer Relevancy metric measures the relevance of the output. Here, GPT-turbo-3.5 is used to extract statements within the output to determine if they are relevant to the input. A higher $\uparrow$ score indicates better results. The rationale for using these evaluation metrics is to gauge bias and toxicity following safety interventions while ensuring the retention of the original content. Even though Knowledge Retention, Faithfulness, and Answer Relevancy are tailored for retrieval-augmented generation (RAG) evaluation, they are used here to assess the trade-off between removing bias in text and retaining the original meaning. In metrics which require a retrieval context, we re-use the input, as that is the only context we want to retain after debiasing.
@@
    Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf: Both the vanilla Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf models were used, each provided with two-shot demonstrations comprising an unsafe example with a neutral variation, to demonstrate safe behavior.
@@
    Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf (both prompt-tuned): The vanilla versions were enhanced with a minimal set of prompt parameters and examples.
These methods were compared against our MBIAS model, which is built on the pre-trained Mistral-7B-Instruct-v0.2 base model using our specialized dataset tailored to this safety task.
@@ 5.1 Overall Results
The analysis in Table 3 explores the comparative performance of different LLMs in terms of bias, toxicity, knowledge retention, faithfulness, and answer relevancy. The table shows results both before and after safety interventions, highlighting the effectiveness of these interventions in reducing bias and toxicity.
@@
Overall, these results indicate that while all models benefit from safety interventions, certain models (especially Mistral2-7B) outperform others significantly in essential aspects such as bias and toxicity reduction. It also suggests that further tuning, especially prompt-tuning, can effectively optimize these models for specific performance metrics.

### v3

@@ Abstract
The deployment of Large Language Models (LLMs) in diverse applications necessitates an assurance of safety without compromising the contextual integrity of the generated content. Traditional approaches, including safety-specific fine-tuning or adversarial testing, often yield safe outputs at the expense of contextual meaning. This can result in a diminished capacity to handle nuanced aspects of bias and toxicity, such as underrepresentation or negative portrayals across various demographics. To address these challenges, we introduce MBIAS, an LLM framework carefully instruction fine-tuned on a custom dataset designed specifically for safety interventions. MBIAS is designed to significantly reduce biases and toxic elements in LLM outputs while preserving the main information. This work also details our further use of LLMs: as annotator under human supervision and as evaluator of generated content. Empirical analysis reveals that MBIAS achieves a reduction in bias and toxicity by over 30% in standard evaluations, and by more than 90% in diverse demographic tests, highlighting the robustness of our approach. We make the dataset and the fine-tuned MBIAS model available to the research community for further investigation and ensure reproducibility. The code for this project can be accessed here  [^1] https://github.com/shainarazavi/MBIAS.
@@
Shaina Raza1,*, Anaya Raval1, Veronica Chatrath1 1Vector Institute, Toronto, ON, Canada. {shaina.raza, ananya.raval, veronica.chatrath} @vectorinstitute.ai
@@ 1 Introduction
The rise of generative artificial intelligence (AI) has brought forth substantial ethical and safety challenges, raising concerns of misuse, misinformation Raza and Ding (2022), and bias of the generated information Wach et al. (2023). Recent initiatives in this line of research for safety in LLMs aim at aligning large language models (LLMs) with ethical norms. These efforts prioritize mitigating biases and enhancing values such as inclusivity, fairness, and non-discrimination Guo et al. (2023). To address harmful, biased, or misleading content (referred to as ‘biased texts’ herein), foundational strategies include implementing safety guardrails in the outputs generated by LLMs. These guardrails provide guidelines and boundaries to ensure AI applications align with ethical standards and societal expectations Att (2024).
@@
Methods such as red-teaming Perez et al. (2022), human feedback during pre-training Casper et al. (2023), and data augmentation methods (e.g., balanced sampling, paraphrasing, or counterfactual data generation) Sadasivan et al. (2023) are often used to reduce biases in LLMs, with the goal of making them safer and more aligned with human values Ganguli et al. (2022); Korbak et al. (2023). In the fine-tuning phase, techniques like instruction tuning, reinforcement learning from human feedback (RLHF), and safety context distillation Ouyang et al. (2022); Qi et al. (2023); Bai et al. (2022) are also used to address unsafe behaviors and improve the models’ ability to generalize.
@@
Despite advancements in the implementation of LLM safety measures, one concern is the loss of actual context or meaning in the original text Schlicht et al. (2024). This delicate balance between two competing goals — reducing biases in the text and preserving informational integrity Raza and Ding (2020) — highlights a central paradox in bias reduction efforts. Catastrophic forgetting, which occurs when a model forgets previously learned information while acquiring new knowledge, is particularly an issue worth noting while implementing safety mechanisms in these LLMs Luo et al. (2023). Demonstrating the understanding of language models post-safety interventions has thus become a topic of research and discussion Nadeem et al. (2021); Schick et al. (2021).
@@
Recent studies indicate that incorporating even a modest number of safety-focused examples during the instruction-tuning stage can effectively mitigate certain safety concerns Bianchi et al. (2023); Inan et al. (2023); Bai et al. (2022). State-of-the-art LLMs such as GPT-4 OpenAI (2023), Mistral Jiang et al. (2023), PaLM Anil et al. (2023), LLaMA-2 Inan et al. (2023) and 3 AI@Meta (2024), and Claude Anthropic (2024) have been further fine-tuned using high-quality safety demonstrations, including perturbations and adversarial prompts, to enhance safety. While not entirely foolproof, this safety-tuning enables LLMs to reduce biases in their outputs. Advancing beyond mere demonstrations, we propose instruction fine-tuning a LLM on safety mechanisms. This approach aligns with prior research that emphasizes the importance of the quality and breadth of instruction-tuning data for developing proficient and reliable instruction-following models Touvron et al. (2023); Wang et al. (2024).
@@
  • We curated an instruction-tuning dataset with a focus on safety considerations. This dataset comprises paired examples: one containing potentially unsafe elements, such as stereotypes or prejudices, and then we present its corresponding benign (safe) counterpart. The dataset was carefully annotated by human experts for reliable ground truth labels (safe counterparts).
  • We propose MBIAS (Mitigating Bias through LLM), an instruction-fine tuned model, built on the top of Mistral2-7B-instruct Jiang et al. (2023). The fine-tuning process involves training the model with our custom dataset that contains examples of both unsafe and safe instances, guiding the model to recognize biases and generate safe responses that can also retain the meaning of the original text.
  • MBIAS can be adapted for use with other LLMs. We utilize parameter-efficient fine-tuning Ding et al. (2023) to train the model, making implementation simple and straightforward. To enhance user-friendliness, we release the model weights, similar to Llama Guard Inan et al. (2023), starting with the smallest available model. This approach ensures that researchers can easily integrate MBIAS into their own projects and benefit from its bias reduction capabilities.
  • In this study, we further explore the versatile roles of LLM as both annotator and judge/ evaluator using GPT-4. Initially, GPT-4 generates accurate ground truth labels for each unsafe input, under human oversight. Later, we employ it as an evaluator, alongside human evaluation, to assess the results of our MBIAS model.
@@ 2 Related Works
Safety in LLMs Establishing safety measures and protocols is crucial to upholding trust in generative AI and LLMs. Many methods have been proposed to address specific biases (that are threat to model safe outputs) in language models. Traditional methods to ensure safe outputs includes examining the embedding space of words and sentences to mitigate biases in the text. Embedding-space-based methods are often applied after training, requiring little-to-no fine-tuning of the model. These methods function as post-processing debiasing steps Liang et al. (2020); Ungless et al. (2022). Subtraction-based methods are also used to remove biases in language models, which map the embedding space into a neutral one Bolukbasi et al. (2016), maintaining equal distance between non-gendered words and pairs of gendered words. Another method is to ensure that gender-neutral words are orthogonal to the gender direction Zhao et al. (2018). In a related work Zhao et al. (2019), the gendered words are replaced with their opposites in the original training data, and the model is trained on both the original and augmented data. Other methods include efficient fine-tuning for debiasing Gira et al. (2022) and fine-tuning only a small portion of parameters on debiasing data Gira et al. (2022).
@@
Debiasing practicality and reliability also depend on dataset selection, evaluation methods, and demographic coverage. Datasets like RedditBias Barikeri et al. (2021), WinoBias Zhao et al. (2019), HolisticBias Smith et al. (2022), RealToxicityPrompts Gehman et al. (2020), and others discussed in Chang et al. (2024) are frequently used for evaluating bias reduction in models. A variety of bias evaluation metrics are used, such as WEAT Liang et al. (2020), Perspective API API (2024), StereoSet Stereotype Score Nadeem et al. (2021), and LLM alignment methods Wang et al. (2023); Chang et al. (2024).
@@
Numerous considerations on their adoption and integration into production settings are currently under observation. Our focus lies in safety implications for instruction-tuned models: specifically, how these models respond safely to biased instructions, while retaining context. Recent research Qian et al. (2022) indicates that training LLMs on demographically perturbed data results in fairer models. We investigate whether training on unsafe-benign text pairs can achieve better fairness in LLMs while retaining their knowledge.
LLM as Annotator Recent advances have showcased LLMs like GPT-3.5 and GPT-4 as promising alternative annotators Tan et al. (2024). These models are capable of annotating data for tasks such as classification and entity recognition AI (2024b), through prompting methods. To maximize the utility of LLMs and leverage vast amounts of unlabeled data, they can be deployed as annotators within an active learning loop Zhang et al. (2023). Furthermore, LLMs annotations can undergo alignment tuning to align their outputs with human preferences Zhao et al. (2023), ensuring the annotations remain free of biases.
LLM based annotations are also shown to achieve or even exceed the performance of human annotators across various NLP tasks, including tweet annotation Huang et al. (2023), computational science Ziems et al. (2024), and medical information extraction Goel et al. (2023). Furthermore, several comparisons between LLMs and human annotators highlight their potential Gilardi et al. (2023); He et al. (2023); Pavlovic and Poesio (2024). Motivated by these findings, our work also employ the role of LLMs in producing benign variations for the biased texts, our goal is to mitigate bias within linguistic contexts while retaining the knowledge —an area that remains relatively unexplored.
LLM as Judge LLMs such as GPT-3.5 and GPT-4 are increasingly utilized as evaluators or judges to ensure outputs align with human preferences Zheng et al. (2024). In this capacity, LLMs can assess and judge the quality of model outputs against specific criteria such as accuracy, toxicity, and relevance Dubois et al. (2024); Zhou et al. (2024), utilizing methods like asking for correctness, relevance or agreement with human annotations in a controlled experiment. In a LLM evaluations  [^2] LMSYS Chatbot Arena (lmsys.org) , OpenAI’s GPT-4 led with 82% of the questions answered correctly, followed by Llama3 and Mixtral. Google’s Gemini Pro 1.5 managed to answer only half of the questions correctly, while GPT-3.5 ranked last. Motivated by these results, we also employ GPT-4 as an LLM judge in this work.
@@ 3.1 Dataset Preparation
The dataset for this study, extracted from our vast collection of about 3.7M records [^3] Newsmediabias data that spans diverse content from news and social media platforms. This dataset is entirely in English and covers many aspects of bias. We carefully chose a subset of 8.5k records for this work (records with various aspects of bias and text lengths exceeding 100), balancing diversity with computational efficiency.
Data Labeling and Annotation Procedure: During labeling, each instance of biased text (unsafe) is transformed into its benign, safe, and debiased counterpart. In the initial step, we leverage GPT-4 from OpenAI OpenAI as a annotator to generate the benign versions of the data. An illustrative prompt (without demonstration) is presented below, while demonstrations are provided in Appendix A:
@@
Please rewrite the following text to create a benign variation that eliminates all forms of bias, toxicity, stereotyping, and harm, such as gender, age, race, political, and socio-economic biases. Ensure the revised text promotes inclusivity and neutrality without sacrificing the original message essence.
@@
Response 1: Leadership skills are not determined by gender. Anyone can be a great leader.
[…]
Response 2: Leadership effectiveness is not limited by age. People of any age can lead well.
[…]
Response 3: Your Turn
@@
The dataset schema [^4] We make our dataset available under the CC BY-NC 4.0 license. is given in Listing 1 below:
@@
To ensure seamless interoperability and integration across computational environments, our dataset follows the Alpaca dataset format Dubois et al. (2023). This structure is comprised of ‘instructions’, ‘input’, and ‘output’.
@@ 3.2 Model Training
To develop MBIAS, we employ QLoRA (Quantized Language Model Optimization with Low Rank Adapters) Belkada et al. (2023), a Parameter-Efficient Fine-tuning (PEFT) technique using bitsandbytes Dettmers et al. (2022), alongside the HuggingFace transformers Trainer class, to fine-tune the Mistral-7B-instruct-0.2555Mistral-7B-Instruct-v0.2 model with our specialized instruction dataset. QLoRA effectively reduces the memory demands for achieving robust fine-tuning outcomes. It balances accuracy with resource efficiency through a 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers. We release our model weights in Huggingface  [^6] MBIAS model.
@@ 4.1 Experimental Setting
Table 2: Hyperparameters used for MBIAS
@@ 4.2 Evaluation Data, Metrics, and Baselines
To evaluate the level of bias and toxicity before and after implementing safety interventions using our methodology, we utilized LLM-based scoring and qualitative measures. When we use LLM as a judge/ evaluator, we use these scoring metrics used in this study, through DeepEval AI (2024a), are defined as follows:
@@
The Knowledge Retention metric measures whether the LLM retains factual information from the input in its generated output. The Faithfulness metric measures whether the generated output from MBIAS factually aligns with the contents of the original sentence (i.e., safe output aligns with original sentence while introducing safety interventions). The Answer Relevancy metric measures the relevance of the output. In this work, GPT-turbo-4 is used to extract statements within the output to determine if they are relevant to the input. A higher $\uparrow$ score indicates better results.
The rationale for using these evaluation metrics is to measure bias and toxicity following safety interventions while ensuring the retention of the original content. Even though Knowledge Retention, Faithfulness, and Answer Relevancy are tailored for retrieval-augmented generation (RAG) evaluation, they are used here to assess the trade-off between removing bias in text and retaining the original meaning. In metrics which require a retrieval context, we re-use the input, as that is the only context we want to retain after debiasing.
@@
    Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf: Both the vanilla Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf models were used using inference, each provided with two-shot demonstrations comprising an unsafe example with a neutral variation, to demonstrate safe behavior.
@@
    Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf (both prompt-tuned): The vanilla versions were enhanced with a minimal set of prompt parameters and examples. Prompt-tuning involves designing specific input prompts (with 2 demonstrations) and providing examples to guide the models towards desired behavior.
Prompt-Tuning involves modifying input prompts to guide model behavior without changing weights, whereas Fine-Tuning adjusts model weights through training on specific datasets. These methods were compared against our MBIAS model.
@@ 5.1 Overall Results
The analysis in Table 3 explores the comparative performance of different LLMs in terms of bias, toxicity, knowledge retention, faithfulness, and answer relevancy. The results show both before and after safety interventions, highlighting the effectiveness of these interventions in reducing bias and toxicity.
@@
Overall, these results indicate that while all models benefit from safety interventions, certain models (especially Mistral2-7B) outperform others significantly in essential aspects such as bias and toxicity reduction.
Main Finding: Fine-tuning LLMs can reduce bias and toxicity while retaining knowledge, faithfulness, and relevance. Prompt-tuning can also serve this purpose, especially when used with already safety fine-tuned models (such as Llama and Mistral models which are already fine-tuned for safety from the providers). However, this approach may result in some knowledge loss but requires less computational resources.